OpenStudy (anonymous):

help!!!i need help on these questions with (11-7) section

OpenStudy (anonymous):

one minute

OpenStudy (anonymous):

i have to do 11-7

OpenStudy (anonymous):

it is difficult to do 11-7

OpenStudy (anonymous):

i don't know what to use for each sine,cosine,tangent,and the equation.Just tell me that.

OpenStudy (anonymous):

Can you use a calculator?

OpenStudy (anonymous):

i would show my work after checking on calculator.I can do that.

OpenStudy (anonymous):

why you don't ask your teacher in the class?

OpenStudy (anonymous):

he doesn't have time today

OpenStudy (anonymous):

nor tomorrrow

OpenStudy (anonymous):

i need to get an A!!!!

OpenStudy (anonymous):

all answers correct

OpenStudy (anonymous):

but it's his obligation to explain to you all you need.

OpenStudy (anonymous):

we need to solve on our own he says..i need to finish it..help me!!!

OpenStudy (anonymous):

i also have to do my other homeworks.

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

well then he ran away!!ha!!!

OpenStudy (zzr0ck3r):

11-7 = 4

OpenStudy (anonymous):

not 11-7 11 dash 4 section to attach file

OpenStudy (zzr0ck3r):

i was joking

OpenStudy (zzr0ck3r):

for the first one, use the law of sines \(\frac{sin(87)}{x}=\frac{3}{.75}\)solce for x

OpenStudy (anonymous):

ok i really need help

OpenStudy (zzr0ck3r):

look up

OpenStudy (anonymous):

i have another homework which will take 2 hours so i can't waste my time

OpenStudy (zzr0ck3r):

have you heard of law of sines?

OpenStudy (zzr0ck3r):

well if its homework, its not a waste of time. The waste of time was the stuff you were doing when you should have been doing your homework

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

well i just came home,i neeed help

OpenStudy (anonymous):

ran away wasters of time!!!

OpenStudy (zzr0ck3r):

I showed you how to do the first one, then asked you a question. Please answer:)

OpenStudy (anonymous):

i don't understand the first one

OpenStudy (zzr0ck3r):

HAVE YOU HEARD OF THE LAW OF SINES?

OpenStudy (anonymous):

don't understand the writing

OpenStudy (anonymous):

yeah don't understand the cursive writing

OpenStudy (zzr0ck3r):

cursive?

OpenStudy (zzr0ck3r):

so you know law of sines does this make sense? \(\large\frac{\sin(87)}{x}=\frac{3}{.75}\)

OpenStudy (anonymous):

write it clearly

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how is it 87

OpenStudy (zzr0ck3r):

we know one angle is 90, we know another is 3, they must all add to 180, so that leaves one option for the third angle

OpenStudy (zzr0ck3r):

you should be nicer when other people are helping you.

OpenStudy (zzr0ck3r):

it will make life much easier

OpenStudy (anonymous):

u would do .0524=.75/x

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

.0524 is the sine of 3 degrees

OpenStudy (anonymous):

u mean 0.9986/x=3/.75

OpenStudy (anonymous):

hellooooooooooo

OpenStudy (anonymous):

didn't even help me

OpenStudy (anonymous):

wasted my time